OpenStudy (anonymous):

A box has a volume of x^3 - 16x cm^3. Another box has a volume of x^3 - 4x^2 cm^3. What is the ratio of the volume of the first box to that of the second? So I divide, right? x^3 - 16 / x^3 - 4x^2 ? But then I'm stuck...

OpenStudy (anonymous):

* x^3 - 16x / x^3 - 4x^2

OpenStudy (isaiah.feynman):

You need to factor like so..

OpenStudy (anonymous):

ok... so \[\frac{ x(x^2- 16) }{ x^2(x -4) }\] \[\frac{ x(x+4)(x-4) }{ x^2(x-4) }\] am I on the right track?

OpenStudy (isaiah.feynman):

Yes you are.

OpenStudy (anonymous):

what do I do next?

OpenStudy (isaiah.feynman):

Find any common factors?

OpenStudy (anonymous):

(x-4)?

OpenStudy (isaiah.feynman):

Yes, and one more.

OpenStudy (anonymous):

x?

OpenStudy (isaiah.feynman):

Excellent!

OpenStudy (anonymous):

so if I take out the common factors... \[\frac{ x + 4 }{ x }\] is that my answer?

OpenStudy (isaiah.feynman):

Yep.

OpenStudy (isaiah.feynman):